OpenStudy (codyhelton):

Match each pair of equivalent fractions. 1.-4a over 8ab 2.-9b^3 over -3a^2b 3.3a^4 over1a^3b^2 4.5a^4b over -10a^2 a) 3b²/1a² b) -1a²b/2 c) -1/2b d) 3a/b²

OpenStudy (mariamikayla):

Where are the pairs?

OpenStudy (codyhelton):

@Mariamikayla

OpenStudy (mariamikayla):

1) -4a/8ab= -a/2ab= -1/2b ---> c) 2) -9b^3/-3a^2b= 3b^3/a^2b= 3b^2/a^2 ---> a) 3) 3a^4/1a^3b^2= 3a/b^2 ---> d) 4) 5a^4b/-10a^2= a^4b/ -2a^2= - a^2b/2 ---> b)